# Beyond the Average: Distributional Causal Inference under Imperfect Compliance

### v1

@@ 1 Introduction
The remainder of the paper is structured as follows. Section 2 reviews related literature. Section 3 describes the problem setup and identification strategy. Section 4 introduces the proposed estimation method. Section 5 presents the asymptotic properties of our estimator. Section 6 reports simulation and empirical results. Section 7 concludes. The techincal appendix (separate file) includes notation, technical proofs, and additional experimental results.
@@ 2 Related Literature
Distributional and quantile treatment effects provide a more comprehensive view of treatment impacts beyond average effects. The concept of QTE was first introduced by Doksum, (1974) and Lehmann and D’Abrera, (1975), and has since inspired a broad literature developing estimation and inference methods for distributional effects across econometrics, statistics, and machine learning. Notable contributions include Heckman et al., (1997); Imbens and Rubin, (1997); Koenker, (2005); Bitler et al., (2006); Athey and Imbens, (2006); Firpo, (2007); Chernozhukov et al., (2013); Koenker et al., (2017); Belloni et al., (2017); Callaway et al., (2018); Callaway and Li, (2019); Chernozhukov et al., (2019); Ge et al., (2020); Park et al., (2021); Zhou et al., (2022); Gunsilius, (2023); Kallus and Oprescu, (2023), among others. Most of this work focuses on conditional distributional and quantile treatment effects. In contrast, Oka et al., (2024) and Byambadalai et al., (2024) study unconditional distributional effects, but under simple random sampling and full compliance.
@@ 3 Setup and Notation
$Y=D\cdot Y(1)+(1-D)\cdot Y(0)\quad\mathrm{and}\quad D=Z\cdot D(1)+(1-Z)\cdot D(0).$
@@
Assumption 3.1 (i) allows for cross-sectional dependence among treatment statuses $\{Z_{i}\}_{i=1}^{n}$, thereby accomodating many covariate-adaptive randomization schemes. Assumption 3.1 (ii) states that the assignment is independent of potential outcomes, potential treatment choices and pre-treatment covariates conditional on strata. Assumption 3.1 (iii) states the assignment probabilities converge to the target assignment probabilities as sample size increases. Common randomization schemes satisfying Assumption 3.1 (i) to (iii) include simple random sampling, stratified block randomization, biased-coin design Efron, (1971), and adaptive biased-coin design Wei, (1978). Assumption 3.1 (iv) says that there are no defiers in the population. This assumption is also called the monotonicity assumption in the literature, and is the key assumption that allows for the identification of the causal effect within a specific subpopulation, known as compliers.
@@
$\mathbb{F}_{Y(d)}(y):=\mathbb{P}\big(Y(d)\leq y\big)\text{ for }d\in\{0,1\},y\in\mathcal{Y}.$
@@
$\beta(y):=$ $\mathbb{F}_{Y(1)}\big(y\mid D(1)>D(0)\big)-\mathbb{F}_{Y(0)}\big(y\mid D(1)>D(0)\big),$
@@ 4 Estimation
$\mu_{z}(y,s,x):=\mathbb{E}[1\!{\rm l}_{\{Y\leq y\}}\mid Z=z,S=s,X=x]\text{ for }y\in\mathcal{Y}.$
@@
$\Xi^{Y}_{z,i}(y)=$ $\frac{1\!{\rm l}_{\{Z_{i}=z\}}\cdot\big(1\!{\rm l}_{\{Y_{i}\leq y\}}-\widehat{\mu}_{z}(y,S_{i},X_{i})\big)}{\widehat{\pi}_{z}(S_{i})}+\widehat{\mu}_{z}(y,S_{i},X_{i}),$
$\Xi^{D}_{z,i}=$ $\frac{1\!{\rm l}_{\{Z_{i}=z\}}\cdot\big(D_{i}-\widehat{\eta}_{z}(S_{i},X_{i})\big)}{\widehat{\pi}_{z}(S_{i})}+\widehat{\eta}_{z}(S_{i},X_{i}),\ \ \ \mathrm{for}\ \ z=0,1.$
@@ 5 Asymptotic Properties
$\sup_{y\in\mathcal{Y},s\in\mathcal{S}}$ $\biggl|\frac{\sum_{i\in I_{1}(s)}\delta^{Y}_{z}(y,s,X_{i})}{n_{1}(s)}-\frac{\sum_{i\in I_{0}(s)}\delta^{Y}_{z}(y,s,X_{i})}{n_{0}(s)}\biggr|=o_{p}(n^{-1/2}),$
$\max_{s\in\mathcal{S}}$ $\biggl|\frac{\sum_{i\in I_{1}(s)}\delta^{D}_{z}(s,X_{i})}{n_{1}(s)}-\frac{\sum_{i\in I_{0}(s)}\delta^{D}_{z}(s,X_{i})}{n_{0}(s)}\biggr|=o_{p}(n^{-1/2}).$
@@
$\sqrt{n}\big(\widehat{\beta}(y)-\beta(y)\big)\rightsquigarrow\mathcal{G}(y),$
@@
$\Omega(y,y^{\prime}):=\frac{\Omega_{0}(y,y^{\prime})+\Omega_{1}(y,y^{\prime})+\Omega_{2}(y,y^{\prime})}{\mathbb{E}[D(1)-D(0)]^{2}},$
@@
    $\Omega(y):=\frac{\Omega_{0}(y,y)+\Omega_{1}(y,y)+\Omega_{2}(y,y)}{\mathbb{E}[D(1)-D(0)]^{2}},$
@@ 6.1 Simulation Study
Figure 2: RMSE, Average Confidence Interval (CI) length, and Coverage Probability (n=1000).
@@
$Y_{i}(d)$ $=a_{d}+b(X_{i},W_{i})+\epsilon_{i}\ \ \text{ for }d\in\{0,1\}$
$D_{i}(0)$ $=1\!{\rm l}_{\{b_{0}+c(X_{i},W_{i})>c_{0}\epsilon_{i}\}},$
@@
Figure 3: RMSE reduction over unadjusted estimator across varying sample sizes
@@ 6.2 Real Data Analysis: Oregon Health Insurance Experiment
Figure 4: Oregon Health Insurance Experiment: Local Distributional Treatment Effect (LDTE) and Local Probability Treatment Effect (LPTE) of insurance coverage on number of emergency department (ED) visits. The left panels depict the empirical probability estimates, while the right panels present regression-adjusted estimates obtained using gradient boosting with 2-fold cross-fitting. Shaded regions and error bars represent 95% confidence intervals. Sample size: n=17,021n=17{,}021.
Figure 4 displays the distributional and probability treatment effect of insurance coverage on ED visits. We compute the LDTE and Local Probability Treatment Effect (LPTE) for $y\in\{0,1,\dots,15\}$ accounting for the stratified design and imperfect compliance. For regression adjustment, we use gradient boosting with 2-fold cross-fitting, with 28 pre-treatment covariates ($X_{i}$) including various variables regarding past emergency department visits. The full list of covariates can be found in the Appendix.
The top-left panel of Figure 4 displays the empirical LDTE, while the top-right panel presents the regression-adjusted LDTE. Shaded areas represent 95% confidence bands, constructed using 500 bootstrap replications. In this case, regression adjustment reduces standard errors by approximately 10–20%. Similarly, the bottom-left panel shows the empirical LPTE, and the bottom-right panel shows the regression-adjusted LPTE, where standard errors are reduced by about 5.5–10% across the distribution.
The distributional analysis reveals that the probability of having zero emergency department visits decreases by 11 percentage points (pp), with a standard error of 4.9 pp (or 4.5 pp with regression adjustment). Beyond this, the only marginally significant effect is an increase of approximately 2 pp in the probability of having five ED visits, with a standard error of 1 pp. No other statistically significant changes are observed across the rest of the distribution, even after applying regression adjustment.
@@ 7 Conclusion
This work has several limitations. It relies on standard IV assumptions such as monotonicity and the exclusion restriction, and focuses on binary treatments. Performance may vary depending on the quality of nuisance estimation in finite samples. Future research could extend the framework to multi-valued or continuous treatments, relax identifying assumptions, and explore dynamic or longitudinal settings.

### v2

@@ 1 Introduction
The remainder of the paper is structured as follows. Section 2 reviews related literature. Section 3 describes the problem setup and identification strategy. Section 4 introduces the proposed estimation method. Section 5 presents the asymptotic properties of our estimator. Section 6 reports simulation and empirical results. Section 7 concludes. The Appendix includes notation, technical proofs, and additional experimental details.
@@ 2 Related Literature
Distributional and quantile treatment effects provide a more comprehensive view of treatment impacts beyond average effects. The concept of QTE was first introduced by Doksum, (1974) and Lehmann and D’Abrera, (1975), and has since inspired a broad literature developing estimation and inference methods for distributional effects across econometrics, statistics, and machine learning. Notable contributions include Heckman et al., (1997); Imbens and Rubin, (1997); Koenker, (2005); Bitler et al., (2006); Athey and Imbens, (2006); Firpo, (2007); Chernozhukov et al., (2013); Koenker et al., (2017); Belloni et al., (2017); Callaway et al., (2018); Callaway and Li, (2019); Chernozhukov et al., (2019); Ge et al., (2020); Park et al., (2021); Zhou et al., (2022); Gunsilius, (2023); Kallus and Oprescu, (2023), among others. Most of this work focuses on conditional distributional and quantile treatment effects. In contrast, Oka et al., (2025), Byambadalai et al., (2024), and Hirata et al., (2025) examine unconditional distributional effects, though their analyses are restricted to settings with simple random sampling and full compliance. Byambadalai et al., (2025) also examine unconditional distributional effects under covariate-adaptive randomization, but their framework likewise assumes full compliance.
@@ 3 Setup and Notation
$Y=D\cdot Y(1)+(1-D)\cdot Y(0)\quad\mathrm{and}\quad D=Z\cdot D(1)+(1-Z)\cdot D(0).$ (1)
@@
Assumption 3.1 (i) allows for cross-sectional dependence among treatment statuses $\{Z_{i}\}_{i=1}^{n}$, thereby accomodating many covariate-adaptive randomization schemes. Assumption 3.1 (ii) states that the assignment is independent of potential outcomes, potential treatment choices and pre-treatment covariates conditional on strata. Assumption 3.1 (iii) states the assignment probabilities converge to the target assignment probabilities as sample size increases.
Common randomization schemes satisfying Assumption 3.1 (i) to (iii) include simple random sampling, stratified block randomization, biased-coin design (Efron,, 1971), and adaptive biased-coin design (Wei,, 1978). Assumption 3.1 (iv) says that there are no defiers in the population. This assumption is also called the monotonicity assumption in the literature, and is the key assumption that allows for the identification of the causal effect within a specific subpopulation, known as compliers.
@@
$\mathbb{F}_{Y(d)}(y):=\mathbb{P}\big(Y(d)\leq y\big)\text{ for }d\in\{0,1\},y\in\mathcal{Y}.$ (2)
@@
$\beta(y):=$ $\mathbb{F}_{Y(1)}\big(y\mid D(1)>D(0)\big)-\mathbb{F}_{Y(0)}\big(y\mid D(1)>D(0)\big),$ (3)
@@
Lastly, we also define the local probability treatment effect (LPTE)
$LPTE(y_{j}):=\mathbb{P}\big(y_{j-1}<Y(1)\leq y_{j}\mid D(1)>D(0)\big)-\mathbb{P}\big(y_{j-1}<Y(0)\leq y_{j}\mid D(1)>D(0)\big),$
for each $j=1,\dots,J$, where $\mathcal{Y}_{J}:=\{y_{1},\cdots,y_{J}\}\subset\mathcal{Y}$ and $y_{0}=$-$\infty$. The LPTE measures treatment-induced changes in the probability mass of the outcome distribution within each interval $(y_{j-1},y_{j}]$, effectively comparing the “histograms” of potential outcomes for compliers. The theoretical results developed for the LDTE extend directly to the LPTE by substituting the indicator functions $1\!{\rm l}_{\{Y(d)\leq y_{j}\}}$ with $1\!{\rm l}_{\{y_{j-1}<Y(d)\leq y_{j}\}}$ for $d\in\{0,1\}$ in all relevant expressions.
@@ 4 Estimation
$\mu_{z}(y,s,x):=\mathbb{E}[1\!{\rm l}_{\{Y\leq y\}}\mid Z=z,S=s,X=x]\text{ for }y\in\mathcal{Y}.$ (6)
@@
$\Xi^{Y}_{z,i}(y)=$ $\frac{1\!{\rm l}_{\{Z_{i}=z\}}\cdot\big(1\!{\rm l}_{\{Y_{i}\leq y\}}-\widehat{\mu}_{z}(y,S_{i},X_{i})\big)}{\widehat{\pi}_{z}(S_{i})}+\widehat{\mu}_{z}(y,S_{i},X_{i}),$ (8)
$\Xi^{D}_{z,i}=$ $\frac{1\!{\rm l}_{\{Z_{i}=z\}}\cdot\big(D_{i}-\widehat{\eta}_{z}(S_{i},X_{i})\big)}{\widehat{\pi}_{z}(S_{i})}+\widehat{\eta}_{z}(S_{i},X_{i}),\ \ \ \mathrm{for}\ \ z=0,1.$ (9)
@@ 5 Asymptotic Properties
$\sup_{y\in\mathcal{Y},s\in\mathcal{S}}$ $\biggl|\frac{\sum_{i\in I_{1}(s)}\delta^{Y}_{z}(y,s,X_{i})}{n_{1}(s)}-\frac{\sum_{i\in I_{0}(s)}\delta^{Y}_{z}(y,s,X_{i})}{n_{0}(s)}\biggr|=o_{p}(n^{-1/2}),$ (10)
$\max_{s\in\mathcal{S}}$ $\biggl|\frac{\sum_{i\in I_{1}(s)}\delta^{D}_{z}(s,X_{i})}{n_{1}(s)}-\frac{\sum_{i\in I_{0}(s)}\delta^{D}_{z}(s,X_{i})}{n_{0}(s)}\biggr|=o_{p}(n^{-1/2}).$ (11)
@@
$\sqrt{n}\big(\widehat{\beta}(y)-\beta(y)\big)\rightsquigarrow\mathcal{G}(y),$ (15)
@@
$\Omega(y,y^{\prime}):=\frac{\Omega_{0}(y,y^{\prime})+\Omega_{1}(y,y^{\prime})+\Omega_{2}(y,y^{\prime})}{\mathbb{E}[D(1)-D(0)]^{2}},$ (16)
@@
    $\Omega(y):=\frac{\Omega_{0}(y,y)+\Omega_{1}(y,y)+\Omega_{2}(y,y)}{\mathbb{E}[D(1)-D(0)]^{2}},$ (17)
@@ 6.1 Simulation Study
Figure 2: Simulation results for LDTE estimators under a nonlinear, high-dimensional design ($n=1000$). RMSE, average 95% CI length, and empirical coverage are shown across quantiles $\{0.1,\dots,0.9\}$ based on 1000 replications. Three estimators are compared: unadjusted, linearly adjusted, and ML-adjusted (gradient boosting with 2-fold cross-fitting). Both adjusted estimators improve RMSE and CI length; the unadjusted estimator attains near-nominal coverage, while ML adjustment slightly over-covers, reflecting conservative inference.
@@
$Y_{i}(d)$ $=a_{d}+b(X_{i},W_{i})+\epsilon_{i}\ \ \text{ for }d\in\{0,1\}$ (18)
$D_{i}(0)$ $=1\!{\rm l}_{\{b_{0}+c(X_{i},W_{i})>c_{0}\epsilon_{i}\}},$ (19)
@@
Figure 3: RMSE reduction (%) of adjusted estimators relative to the unadjusted estimator across quantiles and sample sizes. Linear adjustment yields modest efficiency gains (1–10%), while ML adjustment achieves up to 50% reduction, with improvements becoming more pronounced as sample size increases.
@@ 6.2 Real Data Analysis: Oregon Health Insurance Experiment
Figure 4: Oregon Health Insurance Experiment: Local Distributional Treatment Effect (LDTE) and Local Probability Treatment Effect (LPTE) of insurance coverage on number of emergency department (ED) visits. The left panels depict the empirical probability estimates, while the right panels present regression-adjusted estimates obtained using gradient boosting with 5-fold cross-fitting. Shaded regions and error bars represent 95% confidence intervals. Sample size: n=17,021n=17{,}021.
Figure 4 displays the distributional and probability treatment effect of insurance coverage on ED visits. We compute the LDTE and Local Probability Treatment Effect (LPTE) for $y\in\{0,1,\dots,15\}$ accounting for the stratified design and imperfect compliance. For regression adjustment, we use gradient boosting with 5-fold cross-fitting, with 28 pre-treatment covariates ($X_{i}$) including various variables regarding past emergency department visits. The full list of covariates can be found in the Appendix.
The top-left panel of Figure 4 displays the empirical LDTE, while the top-right panel presents the regression-adjusted LDTE. Shaded areas represent 95% confidence bands, constructed using 500 bootstrap replications. In this case, regression adjustment reduces standard errors by approximately 0.5–15%. Similarly, the bottom-left panel shows the empirical LPTE, and the bottom-right panel shows the regression-adjusted LPTE. Here, the standard errors decrease by about 3.5–26.5% across most of the distribution, except at $y\in\{0,1,2,3\}$, where a slight increase in standard errors is observed.
The regression-adjusted distributional analysis reveals that the probability of having zero emergency department visits decreases by 9 percentage points (pp), with a standard error of 4.2 pp. Beyond this, the only marginally significant effect at the 5% level is an increase of approximately 1.7 pp in the probability of having five ED visits, with a standard error of 0.8 pp. No other statistically significant changes are observed across the rest of the distribution, even after applying regression adjustment.
@@ 7 Conclusion
This work has several limitations. It relies on standard IV assumptions such as monotonicity and the exclusion restriction, and focuses on binary treatments. Performance may vary depending on the quality of nuisance estimation in finite samples. Future research could extend the framework to multi-valued or continuous treatments, relax identifying assumptions, and explore dynamic or longitudinal settings. Furthermore, extending the non-asymptotic frameworks developed by Su et al., 2023a ; Su et al., 2023b to a distributional setting represents a promising avenue for future research.
8 Acknowledgements
We are deeply grateful to the four anonymous reviewers and the program chairs for their thoughtful feedback and constructive discussions, which greatly improved the quality of this paper. Oka also acknowledges the financial support provided by JSPS KAKENHI (Grant Number 24K04821).
NeurIPS Paper Checklist
  1. 1.
    Claims
  2. Question: Do the main claims made in the abstract and introduction accurately reflect the paper’s contributions and scope?
  3. Answer: [Yes]
  4. Justification: The abstract and introduction clearly state the paper’s objectives, proposed approach, and key findings, which are consistently developed and supported throughout the paper.
  5. Guidelines:
    • The answer NA means that the abstract and introduction do not include the claims made in the paper.
    • The abstract and/or introduction should clearly state the claims made, including the contributions made in the paper and important assumptions and limitations. A No or NA answer to this question will not be perceived well by the reviewers.
    • The claims made should match theoretical and experimental results, and reflect how much the results can be expected to generalize to other settings.
    • It is fine to include aspirational goals as motivation as long as it is clear that these goals are not attained by the paper.
  6. 2.
    Limitations
  7. Question: Does the paper discuss the limitations of the work performed by the authors?
  8. Answer: [Yes]
  9. Justification: The paper discusses the limitations of the work in the Conclusion section and suggests directions for future research.
  10. Guidelines:
    • The answer NA means that the paper has no limitation while the answer No means that the paper has limitations, but those are not discussed in the paper.
    • The authors are encouraged to create a separate "Limitations" section in their paper.
    • The paper should point out any strong assumptions and how robust the results are to violations of these assumptions (e.g., independence assumptions, noiseless settings, model well-specification, asymptotic approximations only holding locally). The authors should reflect on how these assumptions might be violated in practice and what the implications would be.
    • The authors should reflect on the scope of the claims made, e.g., if the approach was only tested on a few datasets or with a few runs. In general, empirical results often depend on implicit assumptions, which should be articulated.
    • The authors should reflect on the factors that influence the performance of the approach. For example, a facial recognition algorithm may perform poorly when image resolution is low or images are taken in low lighting. Or a speech-to-text system might not be used reliably to provide closed captions for online lectures because it fails to handle technical jargon.
    • The authors should discuss the computational efficiency of the proposed algorithms and how they scale with dataset size.
    • If applicable, the authors should discuss possible limitations of their approach to address problems of privacy and fairness.
    • While the authors might fear that complete honesty about limitations might be used by reviewers as grounds for rejection, a worse outcome might be that reviewers discover limitations that aren’t acknowledged in the paper. The authors should use their best judgment and recognize that individual actions in favor of transparency play an important role in developing norms that preserve the integrity of the community. Reviewers will be specifically instructed to not penalize honesty concerning limitations.
  11. 3.
    Theory assumptions and proofs
  12. Question: For each theoretical result, does the paper provide the full set of assumptions and a complete (and correct) proof?
  13. Answer: [Yes]
  14. Justification: The paper provides a complete and correct set of assumptions for each theoretical result, with all theorems and proofs clearly stated and appropriately referenced. Full proofs are included in the supplemental material, with some explanation in the main text to aid understanding.
  15. Guidelines:
    • The answer NA means that the paper does not include theoretical results.
    • All the theorems, formulas, and proofs in the paper should be numbered and cross-referenced.
    • All assumptions should be clearly stated or referenced in the statement of any theorems.
    • The proofs can either appear in the main paper or the supplemental material, but if they appear in the supplemental material, the authors are encouraged to provide a short proof sketch to provide intuition.
    • Inversely, any informal proof provided in the core of the paper should be complemented by formal proofs provided in appendix or supplemental material.
    • Theorems and Lemmas that the proof relies upon should be properly referenced.
  16. 4.
    Experimental result reproducibility
  17. Question: Does the paper fully disclose all the information needed to reproduce the main experimental results of the paper to the extent that it affects the main claims and/or conclusions of the paper (regardless of whether the code and data are provided or not)?
  18. Answer: [Yes]
  19. Justification: The paper proposes a new algorithm and provides all necessary details to reproduce the main experimental results, including a clear description of the algorithm, experimental setup, hyperparameters, and evaluation protocols.
  20. Guidelines:
    • The answer NA means that the paper does not include experiments.
    • If the paper includes experiments, a No answer to this question will not be perceived well by the reviewers: Making the paper reproducible is important, regardless of whether the code and data are provided or not.
    • If the contribution is a dataset and/or model, the authors should describe the steps taken to make their results reproducible or verifiable.
    • Depending on the contribution, reproducibility can be accomplished in various ways. For example, if the contribution is a novel architecture, describing the architecture fully might suffice, or if the contribution is a specific model and empirical evaluation, it may be necessary to either make it possible for others to replicate the model with the same dataset, or provide access to the model. In general. releasing code and data is often one good way to accomplish this, but reproducibility can also be provided via detailed instructions for how to replicate the results, access to a hosted model (e.g., in the case of a large language model), releasing of a model checkpoint, or other means that are appropriate to the research performed.
    • While NeurIPS does not require releasing code, the conference does require all submissions to provide some reasonable avenue for reproducibility, which may depend on the nature of the contribution. For example
      1. (a)
        If the contribution is primarily a new algorithm, the paper should make it clear how to reproduce that algorithm.
      2. (b)
        If the contribution is primarily a new model architecture, the paper should describe the architecture clearly and fully.
      3. (c)
        If the contribution is a new model (e.g., a large language model), then there should either be a way to access this model for reproducing the results or a way to reproduce the model (e.g., with an open-source dataset or instructions for how to construct the dataset).
      4. (d)
        We recognize that reproducibility may be tricky in some cases, in which case authors are welcome to describe the particular way they provide for reproducibility. In the case of closed-source models, it may be that access to the model is limited in some way (e.g., to registered users), but it should be possible for other researchers to have some path to reproducing or verifying the results.
  21. 5.
    Open access to data and code
  22. Question: Does the paper provide open access to the data and code, with sufficient instructions to faithfully reproduce the main experimental results, as described in supplemental material?
  23. Answer: [Yes]
  24. Justification: The paper provides open access to both the code and data (we use publicly available data and we included the link to download), along with detailed instructions in the supplemental material for setting up the environment and reproducing the main experimental results.
  25. Guidelines:
    • The answer NA means that paper does not include experiments requiring code.
    • Please see the NeurIPS code and data submission guidelines (https://nips.cc/public/guides/CodeSubmissionPolicy) for more details.
    • While we encourage the release of code and data, we understand that this might not be possible, so “No” is an acceptable answer. Papers cannot be rejected simply for not including code, unless this is central to the contribution (e.g., for a new open-source benchmark).
    • The instructions should contain the exact command and environment needed to run to reproduce the results. See the NeurIPS code and data submission guidelines (https://nips.cc/public/guides/CodeSubmissionPolicy) for more details.
    • The authors should provide instructions on data access and preparation, including how to access the raw data, preprocessed data, intermediate data, and generated data, etc.
    • The authors should provide scripts to reproduce all experimental results for the new proposed method and baselines. If only a subset of experiments are reproducible, they should state which ones are omitted from the script and why.
    • At submission time, to preserve anonymity, the authors should release anonymized versions (if applicable).
    • Providing as much information as possible in supplemental material (appended to the paper) is recommended, but including URLs to data and code is permitted.
  26. 6.
    Experimental setting/details
  27. Question: Does the paper specify all the training and test details (e.g., data splits, hyperparameters, how they were chosen, type of optimizer, etc.) necessary to understand the results?
  28. Answer: [Yes]
  29. Justification: The paper thoroughly outlines all relevant experimental details, including the use of a data splitting method known as cross-fitting, the selection and tuning of hyperparameters, and other implementation specifics essential for fully understanding and interpreting the experimental results.
  30. Guidelines:
    • The answer NA means that the paper does not include experiments.
    • The experimental setting should be presented in the core of the paper to a level of detail that is necessary to appreciate the results and make sense of them.
    • The full details can be provided either with the code, in appendix, or as supplemental material.
  31. 7.
    Experiment statistical significance
  32. Question: Does the paper report error bars suitably and correctly defined or other appropriate information about the statistical significance of the experiments?
  33. Answer: [Yes]
  34. Justification: The paper reports error bars for the key experimental results, clearly explaining the computation methods. Bootstrap resampling and analytical standard error calculations were employed to estimate variance and assess statistical significance.
  35. Guidelines:
    • The answer NA means that the paper does not include experiments.
    • The authors should answer "Yes" if the results are accompanied by error bars, confidence intervals, or statistical significance tests, at least for the experiments that support the main claims of the paper.
    • The factors of variability that the error bars are capturing should be clearly stated (for example, train/test split, initialization, random drawing of some parameter, or overall run with given experimental conditions).
    • The method for calculating the error bars should be explained (closed form formula, call to a library function, bootstrap, etc.)
    • The assumptions made should be given (e.g., Normally distributed errors).
    • It should be clear whether the error bar is the standard deviation or the standard error of the mean.
    • It is OK to report 1-sigma error bars, but one should state it. The authors should preferably report a 2-sigma error bar than state that they have a 96% CI, if the hypothesis of Normality of errors is not verified.
    • For asymmetric distributions, the authors should be careful not to show in tables or figures symmetric error bars that would yield results that are out of range (e.g. negative error rates).
    • If error bars are reported in tables or plots, The authors should explain in the text how they were calculated and reference the corresponding figures or tables in the text.
  36. 8.
    Experiments compute resources
  37. Question: For each experiment, does the paper provide sufficient information on the computer resources (type of compute workers, memory, time of execution) needed to reproduce the experiments?
  38. Answer: [Yes]
  39. Justification: The paper provides detailed information on the computational resources used for each experiment, including the type of compute, memory specifications, and execution time.
  40. Guidelines:
    • The answer NA means that the paper does not include experiments.
    • The paper should indicate the type of compute workers CPU or GPU, internal cluster, or cloud provider, including relevant memory and storage.
    • The paper should provide the amount of compute required for each of the individual experimental runs as well as estimate the total compute.
    • The paper should disclose whether the full research project required more compute than the experiments reported in the paper (e.g., preliminary or failed experiments that didn’t make it into the paper).
  41. 9.
    Code of ethics
  42. Question: Does the research conducted in the paper conform, in every respect, with the NeurIPS Code of Ethics https://neurips.cc/public/EthicsGuidelines?
  43. Answer: [Yes]
  44. Justification: The research fully conforms to the NeurIPS Code of Ethics. The study considers potential societal and environmental impacts, avoids known risks such as discrimination or misuse, and follows best practices for reproducibility, transparency, and responsible data handling.
  45. Guidelines:
    • The answer NA means that the authors have not reviewed the NeurIPS Code of Ethics.
    • If the authors answer No, they should explain the special circumstances that require a deviation from the Code of Ethics.
    • The authors should make sure to preserve anonymity (e.g., if there is a special consideration due to laws or regulations in their jurisdiction).
  46. 10.
    Broader impacts
  47. Question: Does the paper discuss both potential positive societal impacts and negative societal impacts of the work performed?
  48. Answer: [Yes]
  49. Justification: The paper discusses potential societal impacts, noting that the proposed method can lead to improved decision-making in applied settings as a positive outcome. It also acknowledges a potential negative impact, namely that the underlying assumptions of the method may not hold in all real-world scenarios, which could limit its effectiveness or lead to unintended consequences.
  50. Guidelines:
    • The answer NA means that there is no societal impact of the work performed.
    • If the authors answer NA or No, they should explain why their work has no societal impact or why the paper does not address societal impact.
    • Examples of negative societal impacts include potential malicious or unintended uses (e.g., disinformation, generating fake profiles, surveillance), fairness considerations (e.g., deployment of technologies that could make decisions that unfairly impact specific groups), privacy considerations, and security considerations.
    • The conference expects that many papers will be foundational research and not tied to particular applications, let alone deployments. However, if there is a direct path to any negative applications, the authors should point it out. For example, it is legitimate to point out that an improvement in the quality of generative models could be used to generate deepfakes for disinformation. On the other hand, it is not needed to point out that a generic algorithm for optimizing neural networks could enable people to train models that generate Deepfakes faster.
    • The authors should consider possible harms that could arise when the technology is being used as intended and functioning correctly, harms that could arise when the technology is being used as intended but gives incorrect results, and harms following from (intentional or unintentional) misuse of the technology.
    • If there are negative societal impacts, the authors could also discuss possible mitigation strategies (e.g., gated release of models, providing defenses in addition to attacks, mechanisms for monitoring misuse, mechanisms to monitor how a system learns from feedback over time, improving the efficiency and accessibility of ML).
  51. 11.
    Safeguards
  52. Question: Does the paper describe safeguards that have been put in place for responsible release of data or models that have a high risk for misuse (e.g., pretrained language models, image generators, or scraped datasets)?
  53. Answer: [N/A]
  54. Justification: The paper does not involve the release of models or datasets that pose a high risk for misuse, such as pretrained language models, generative systems, or scraped data, and therefore no specific safeguards are necessary.
  55. Guidelines:
    • The answer NA means that the paper poses no such risks.
    • Released models that have a high risk for misuse or dual-use should be released with necessary safeguards to allow for controlled use of the model, for example by requiring that users adhere to usage guidelines or restrictions to access the model or implementing safety filters.
    • Datasets that have been scraped from the Internet could pose safety risks. The authors should describe how they avoided releasing unsafe images.
    • We recognize that providing effective safeguards is challenging, and many papers do not require this, but we encourage authors to take this into account and make a best faith effort.
  56. 12.
    Licenses for existing assets
  57. Question: Are the creators or original owners of assets (e.g., code, data, models), used in the paper, properly credited and are the license and terms of use explicitly mentioned and properly respected?
  58. Answer: [Yes]
  59. Justification: The Oregon Health Insurance Experiment dataset is publicly available through the NBER website, and we have appropriately cited the original study in our work.
  60. Guidelines:
    • The answer NA means that the paper does not use existing assets.
    • The authors should cite the original paper that produced the code package or dataset.
    • The authors should state which version of the asset is used and, if possible, include a URL.
    • The name of the license (e.g., CC-BY 4.0) should be included for each asset.
    • For scraped data from a particular source (e.g., website), the copyright and terms of service of that source should be provided.
    • If assets are released, the license, copyright information, and terms of use in the package should be provided. For popular datasets, paperswithcode.com/datasets has curated licenses for some datasets. Their licensing guide can help determine the license of a dataset.
    • For existing datasets that are re-packaged, both the original license and the license of the derived asset (if it has changed) should be provided.
    • If this information is not available online, the authors are encouraged to reach out to the asset’s creators.
  61. 13.
    New assets
  62. Question: Are new assets introduced in the paper well documented and is the documentation provided alongside the assets?
  63. Answer: [Yes]
  64. Justification: The new assets introduced in the paper are well documented, with detailed descriptions of their structure, usage, and limitations. All relevant materials are included as an anonymized zip file in the supplemental submission.
  65. Guidelines:
    • The answer NA means that the paper does not release new assets.
    • Researchers should communicate the details of the dataset/code/model as part of their submissions via structured templates. This includes details about training, license, limitations, etc.
    • The paper should discuss whether and how consent was obtained from people whose asset is used.
    • At submission time, remember to anonymize your assets (if applicable). You can either create an anonymized URL or include an anonymized zip file.
  66. 14.
    Crowdsourcing and research with human subjects
  67. Question: For crowdsourcing experiments and research with human subjects, does the paper include the full text of instructions given to participants and screenshots, if applicable, as well as details about compensation (if any)?
  68. Answer: [N/A]
  69. Justification: The paper does not involve crowdsourcing or research with human subjects, and therefore no participant instructions, screenshots, or compensation details are applicable.
  70. Guidelines:
    • The answer NA means that the paper does not involve crowdsourcing nor research with human subjects.
    • Including this information in the supplemental material is fine, but if the main contribution of the paper involves human subjects, then as much detail as possible should be included in the main paper.
    • According to the NeurIPS Code of Ethics, workers involved in data collection, curation, or other labor should be paid at least the minimum wage in the country of the data collector.
  71. 15.
    Institutional review board (IRB) approvals or equivalent for research with human subjects
  72. Question: Does the paper describe potential risks incurred by study participants, whether such risks were disclosed to the subjects, and whether Institutional Review Board (IRB) approvals (or an equivalent approval/review based on the requirements of your country or institution) were obtained?
  73. Answer: [N/A]
  74. Justification: The paper does not involve research with human subjects or crowdsourcing, so there were no participant risks to assess and no need for IRB or equivalent ethical review.
  75. Guidelines:
    • The answer NA means that the paper does not involve crowdsourcing nor research with human subjects.
    • Depending on the country in which research is conducted, IRB approval (or equivalent) may be required for any human subjects research. If you obtained IRB approval, you should clearly state this in the paper.
    • We recognize that the procedures for this may vary significantly between institutions and locations, and we expect authors to adhere to the NeurIPS Code of Ethics and the guidelines for their institution.
    • For initial submissions, do not include any information that would break anonymity (if applicable), such as the institution conducting the review.
  76. 16.
    Declaration of LLM usage
  77. Question: Does the paper describe the usage of LLMs if it is an important, original, or non-standard component of the core methods in this research? Note that if the LLM is used only for writing, editing, or formatting purposes and does not impact the core methodology, scientific rigorousness, or originality of the research, declaration is not required.
  78. Answer: [N/A]
  79. Justification: The core methodology and contributions of the paper do not involve the use of LLMs in any important, original, or non-standard way.
  80. Guidelines:
    • The answer NA means that the core method development in this research does not involve LLMs as any important, original, or non-standard components.
    • Please refer to our LLM policy (https://neurips.cc/Conferences/2025/LLM) for what should or should not be described.